A random variable X has a normal distribution with mean 50 and variance of 10. Find the range of scores that lie:


A.Within one standard deviation from the mean.


B.Within two standard deviation from the mean.



1
Expert's answer
2022-03-09T12:00:31-0500


Mean: "\mu=50;"

variance: "\sigma^2=10;"

standard deviation: "\sigma=\sqrt{10} \approx3.16."


A. The range of scores that lie within one standard deviation from the mean:

"(\mu-\sigma, \mu+\sigma) =(50-3.16, 50+3.16)="

"=(46.84, 53.16)."


B. The range of scores that lie within two standard deviations from the mean:

"(\mu-2\sigma, \mu+2\sigma) =(50-2\cdot3.16, 50+2\cdot3.16)="

"=(43.68, 56.32)."